# Multipoint Schwarz–Pick Lemma for quaternionic case

### 15 Dec 2023 v1

@@ 5.3. Nevanlinna–Pick interpolation with more than three points
$f(r_{m})=s_{m}\quad\text{for all }m=1,\ldots,n,\qquad f\in\mathcal{SR(\mathbb{B},\mathbb{B})}.$
@@
Remark 5.10.
The condition for having an infinite family of solutions (non-singular case) for the Pick interpolation problem is $|Q_{n-1}^{n}|<1$, or equivalently $|Q_{\kappa}^{\ell}|<1$ for all $\kappa=1,\ldots,n-1$ and $\ell=\kappa+1,\ldots,n$. This can be easily restated in terms of the pseudo-hyperbolic distance $\rho_{\mathbb{B}}$, analogously as in Corollary 5.8: the problem has infinite solutions if and only if $\rho_{\mathbb{B}}(Q_{n-2}^{n-1},Q_{n-2}^{n})<\rho_{\mathbb{B}}(r_{n-1},r_{n})$, or equivalently, if $\rho_{\mathbb{B}}(Q_{\kappa-1}^{\kappa},Q_{\kappa-1}^{\ell})<\rho_{\mathbb{B}}(r_{\kappa},r_{\ell})$ for all $\kappa\in\{1,\ldots,n-1\},\ell\in\{\kappa+1,\ldots,n\}$ – where $Q_{0}^{\ell}:=s_{\ell}$.
@@ 5.4. The case of nodes not in the real line
and the last term evaluates as $M_{p_{1}}(\tilde{p}_{\ell})^{-1}M_{s_{1}}(f(\hat{p}_{\ell}))$ where $\tilde{p}_{\ell}$ is a point in $\mathbb{S}_{p_{\ell}}$ depending on $p_{1}$ and $\hat{p}_{\ell}\in\mathbb{S}_{p_{\ell}}$ may depend on $p_{1},s_{1}$ and on the conjugate function $f^{c}$. They trivially coincide with $p_{\ell}$ when that point is real. So, the proposed algorithm does not work anymore, as $\hat{p}_{\ell}$ generally depends on the function $f^{c}$ which is not known a priori and in any case $f(\hat{p}_{\ell})$ is not known if $\hat{p}_{\ell}$ is not one of the nodes. However, there is a specific situation in which we can recover a partial result in the spirit of Theorem 5.9. As Theorem 5.5 holds in the complex setting ([6, Thm. 4.1]) without the restriction to real nodes of interpolation, then we have the following partial answer: if $p_{1},\ldots,p_{n}\in\mathbb{B}$ belong to a common slice $\mathbb{C}_{I}$ and also $s_{1},\ldots,s_{3}\in\mathbb{B}\cap\mathbb{C}_{I}$, then the interpolation problem $f(p_{m})=s_{m}$, $m\in\{1,\ldots,n\}$ admits infinite solutions that are $\mathbb{C}_{I}$-preserving if $|Q_{n-1}^{n}|<1$, while it admits at least a solution given by a regular Blaschke product of degree at most $n-1$, when $|Q_{n-1}^{n}|=1$, and this is the unique one preserving $\mathbb{C}_{I}$.
@@
Fix $J$ as above and let $\lambda=\operatorname{Re}(JI)$. Then $|\lambda|\leq 1$ and there exists $K\in\{I\}^{\perp}\subset\mathbb{S}$ such that $JI=\lambda+\sqrt{1-\lambda^{2}}K$, so

### 25 Jan 2024 v2

@@ 5.3. Nevanlinna–Pick interpolation with more than three points
(28) $f(r_{m})=s_{m}\quad\text{for all }m=1,\ldots,n,\qquad f\in\mathcal{SR(\mathbb{B},\mathbb{B})}.$
@@
The condition for having an infinite family of solutions (non-singular case) for the Pick interpolation problem is $|Q_{n-1}^{n}|<1$, or equivalently $|Q_{\kappa}^{\ell}|<1$ for all $\kappa=1,\ldots,n-1$ and $\ell=\kappa+1,\ldots,n$. This can be easily restated in terms of the pseudo-hyperbolic distance $\rho_{\mathbb{B}}$, analogously as in Corollary 5.8.
Corollary 5.10.
The interpolation problem (28) has infinite solutions if and only if
$\rho_{\mathbb{B}}(Q_{n-2}^{n-1},Q_{n-2}^{n})<\rho_{\mathbb{B}}(r_{n-1},r_{n}),$
or equivalently, if $\rho_{\mathbb{B}}(Q_{\kappa-1}^{\kappa},Q_{\kappa-1}^{\ell})<\rho_{\mathbb{B}}(r_{\kappa},r_{\ell})$ for all $\kappa\in\{1,\ldots,n-1\},\ell\in\{\kappa+1,\ldots,n\}$ – where we use the notation $Q_{0}^{\ell}:=s_{\ell}$.
@@ 5.4. The case of nodes not in the real line
and the last term evaluates as $M_{p_{1}}(\tilde{p}_{\ell})^{-1}M_{s_{1}}(f(\hat{p}_{\ell}))$ where $\tilde{p}_{\ell}$ is a point in $\mathbb{S}_{p_{\ell}}$ depending on $p_{1}$ and $\hat{p}_{\ell}\in\mathbb{S}_{p_{\ell}}$ may depend on $p_{1},s_{1}$ and on the conjugate function $f^{c}$. They trivially coincide with $p_{\ell}$ when that point is real. So, the proposed algorithm does not work anymore, as $\hat{p}_{\ell}$ generally depends on the function $f^{c}$ which is not known a priori and in any case $f(\hat{p}_{\ell})$ is not known if $\hat{p}_{\ell}$ is not one of the nodes. However, there is a specific situation in which we can recover a partial result in the spirit of Theorem 5.9. As Theorem 5.5 holds in the complex setting ([6, Thm. 4.1]) without the restriction to real nodes of interpolation, then we have the following partial answer: if $p_{1},\ldots,p_{n}\in\mathbb{B}$ belong to a common slice $\mathbb{C}_{I}$ and also $s_{1},\ldots,s_{n}\in\mathbb{B}\cap\mathbb{C}_{I}$, then the interpolation problem $f(p_{m})=s_{m}$, $m\in\{1,\ldots,n\}$ admits infinite solutions that are $\mathbb{C}_{I}$-preserving if $|Q_{n-1}^{n}|<1$, while it admits at least a solution given by a regular Blaschke product of degree at most $n-1$, when $|Q_{n-1}^{n}|=1$, and this is the unique one preserving $\mathbb{C}_{I}$.
@@
Fix $J$ as above and let $\lambda=\operatorname{Re}(JI)$. Then $|\lambda|\leq 1$ and there exists $K$ orthogonal to $I$ such that $JI=\lambda+\sqrt{1-\lambda^{2}}K$, so